# Predicting Change, Not States: An Alternate Framework for Neural PDE Surrogates

### v1

@@ Abstract
Neural surrogates for partial differential equations (PDEs) have become popular due to their potential to quickly simulate physics. With a few exceptions, neural surrogates generally treat the forward evolution of time-dependent PDEs as a black box by directly predicting the next state. While this is a natural and easy framework for applying neural surrogates, it can be an over-simplified and rigid framework for predicting physics. In this work, we propose an alternate framework in which neural solvers predict the temporal derivative and an ODE integrator forwards the solution in time, which has little overhead and is broadly applicable across model architectures and PDEs. We find that by simply changing the training target and introducing numerical integration during inference, neural surrogates can gain accuracy and stability. Predicting temporal derivatives also allows models to not be constrained to a specific temporal discretization, allowing for flexible time-stepping during inference or training on higher-resolution PDE data. Lastly, we investigate why this new framework can be beneficial and in what situations does it work well.
@@ 1 Introduction
Solving equations of this form are of great interest, motivating centuries of research and leading to many modern engineering and scientific advances. Although solutions are analytically intractable, solutions can be approximated by discretizing the domain, where $t_{n}\in\{t_{0},\ldots,t_{N}\}$, $x_{m}\in\{x_{0},\ldots,x_{M}\}$, and the solution $\mathbf{u}(t_{n},x_{m})$ is discretized in this domain. Within this setup, initial conditions $\mathbf{u}(0,x_{m})$ and boundary conditions $B[\mathbf{u}](t_{n},x_{b}),x_{b}=\{x_{m}:x_{m}\in\partial\Omega\}$ are usually given based on the practical application of the PDE.
@@
Neural Surrogates As a result, deriving and solving an accurate approximation of $\mathbf{u}$ can require significant technical expertise and computational resources. This has motivated a new class of deep learning methods, or neural surrogates, that can learn this update from data, rather than an analytical derivation, and be quickly queried [3, 4]. This field of neural surrogates is highly diverse and is rapidly growing, with numerous surrogates proposed based on graph neural networks (GNNs) [5, 6], transformers [7, 8], or convolutional neural networks (CNNs) [9, 10]. In addition to advancements in neural network architectures, many works have aimed to improve different aspects of neural surrogates, such as improving prediction accuracies [11], adapting to irregularly discretized data [12, 13], and scaling models across different datasets [14, 15, 16, 17]. The relevant literature is too vast to detail here; however, for a list of representative neural surrogate works and their numerical baselines, we recommend McGreivy and Hakim [18] and for a survey on neural surrogates and operator learning we recommend Kovachki et al. [19].
@@
Contributions Neural surrogates often treat solution updates as a black box and directly predict the next state, which we call state prediction ($\mathbf{u}(t_{n+1})=F_{\theta}(\mathbf{u}(t_{n}))$). State prediction has become the dominant framework for training and querying neural surrogates, as it is the most obvious way to step forward in time, is easy to implement, and allows neural surrogates to take larger steps $\Delta t$. However, we believe that this can be an over-simplification and, inspired by numerical methods, we ask if predicting the temporal derivative or derivative prediction ($\frac{\partial\mathbf{u}}{\partial t}|_{t=t_{n}}=F_{\theta}(\mathbf{u}(t_{n}))$) can improve neural surrogate accuracies. Importantly, derivative prediction changes the training objective and inference procedure, and this new framework does not contribute additional training cost but can introduce additional inference costs. Fortunately, this is the only practical difference between the two frameworks, allowing derivative prediction to be used with any neural surrogate architecture and generally applied to the same range of PDEs, discretizations, and problem setups.
@@
One may ask why it is necessary to investigate alternative neural surrogate training and inference frameworks. We illustrate its differences in Figure 1 and describe its potential benefits here. Our hypothesis is that the main benefit of derivative prediction is that it allows the use of an ODE integrator to evolve the solution based on the predicted temporal derivative, which can be quickly queried at inference time. This effectively decouples the prediction problem, where the model learns the spatial dynamics that create the change in solution values, and the numerical integrator forwards the solution in time based on this learned derivative. In many PDE setups, this can greatly simplify the learning objective; rather than needing to collapse a spatial and temporal update under a single model prediction (state prediction), models can focus on learning just the change in solution values, which tend to be more important than the actual value of the solutions. Through experimentation, we show that this can result in more accurate model predictions.
@@
Secondly, derivative prediction can be more flexible during inference than state prediction. State prediction fixes the temporal resolution of the solution to the dataset resolution, whereas derivative prediction can generate solutions of arbitrary temporal resolutions by modifying $\Delta t$ in the ODE integrator during inference, which could additionally introduce adaptive step sizing for neural surrogates. Furthermore, derivative prediction can take advantage of different integration schemes, such as multi-step, predictor-corrector, or Runge-Kutta methods to improve or adjust solutions during inference. Importantly, under derivative prediction, these modifications during inference do not require re-training the model. We present experimental results demonstrating these additional capabilities.
@@
A subtle but interesting consequence of this is that models trained with derivative prediction can use PDE data more effectively than training in the conventional state prediction framework. A common practice when training neural surrogates is to down-sample PDE data in time, which is usually finely discretized due to numerical solver constraints, and often up to factors of 102-104. This is necessary since predicting the next timestep directly implicitly fixes the temporal resolution of the learned model; training on the full-resolution data would require models to auto-regressively predict a large number of timesteps due to its fine resolution and therefore quickly accumulate error. Unfortunately, downsampling by large factors results in discarding a considerable amount of PDE data, but since derivative prediction does not fix the model to a temporal resolution, this can be avoided. In particular, models can be trained on the full-resolution data to obtain highly accurate estimates of the instantaneous derivative, which can be used to take variable steps in time when numerically integrating during inference. We provide results demonstrating this as well.
@@
The benefits of the derivative prediction are not without drawbacks. The main limitation of derivative prediction is that it reintroduces numerical error and discretization constraints back into the neural surrogate framework. Specifically, even with perfect estimates of the instantaneous derivative, numerical integrators can still accumulate error over time with large timesteps $\Delta t$ or chaotic PDE dynamics. While this is an issue, we find that in practice, this additional numerical error is very small when compared to model error, especially when using higher-order integrators. Additionally, common timescales that conventional neural surrogates are trained on are generally small enough for ODE integrators to handle, however this is not always the case. This can be seen for steady-state problems where neural surrogates are tasked with predicting a future steady-state solution given solely the initial condition (such as Darcy Flow); derivative prediction will completely fail since the initial derivative has little correlation to the steady-state solution. While we only consider time-dependent PDEs, we seek to understand the limits of neural surrogates when predicting different timescales, both under state prediction and derivative prediction frameworks. In particular, it is known that neural surrogates can take timesteps that are numerically instable (e.g., $CFL>1$), but we seek to understand to what extent this is possible. Lastly, we release the code and datasets for this work here: https://github.com/anthonyzhou-1/temporal_pdes.
@@ 2 Related Work
Neural Surrogate Modifications Initial work on neural surrogates designed various architectures to quickly and accurately approximate PDE solutions, including DeepONet [4], neural operators [3], and physics-informed methods [20]. Although powerful on their own, subsequent works have proposed modifications to these architectures to improve accuracy [21, 22, 23, 24], generalize to different problems [14, 8], or adapt to complex PDE systems [25, 26, 27]. While these modifications are insightful and perform well, they are generally problem- or model-specific, which limits their applicability. To address this, prior work has also considered proposing training modifications or guidelines to improve surrogate accuracies [16]. One such example is the use of data augmentations to improve neural surrogate training and prediction accuracy [28]. In addition, previous work has explored the idea of noising data or unrolled training to improve neural surrogate stability and robustness [12, 14, 29]. Lastly, an interesting set of works seek to train refiner modules alongside neural surrogates to improve solution accuracy [11].
@@
Residual Prediction Prior work has also examined learning physics under different training and prediction frameworks. Besides state prediction, the most common training framework is to predict the residual between the input and the target, or residual prediction. Residual prediction can be seen as a special case of derivative prediction, where the residual is a scaled forward Euler approximation of the temporal derivative; however, the temporal resolution is still fixed during inference. Despite residual prediction being used widely in weather applications [30, 31], its use in PDE applications can only be found in certain models [32, 33, 34, 35, 27], with its adoption on an ad hoc basis. Therefore, this paper seeks to understand the benefits/drawbacks of different prediction frameworks to better motivate residual prediction and more broadly, derivative prediction.
@@
Hybrid Solvers Another set of more restrictive, but more targeted training frameworks is to use the neural surrogate to approximate quantities within a conventional solver, creating a hybrid solver. For example, when considering the Navier-Stokes equations, hybrid solvers can use a neural surrogate to approximate the convective flux due to its large computational burden and solve the remaining terms or steps numerically [36, 37]. For other equations, spatial derivatives can also be individually approximated using a neural surrogate [38], and more generally, neural surrogates have found use in approximating costly terms within numerical solvers using data [39, 40]. While these frameworks also simplify the training objective by focusing on a single phenomena to model, they are tied to a specific problem or numerical solver.
@@ 3.1 Training
During training, models parameterized by $F_{\theta}$ are given the current solution $\mathbf{u}(t_{n})$ and timestep $t_{n}$, and are trained to predict either $\mathbf{u}(t_{n+1})$ (state prediction) or $\frac{\partial\mathbf{u}}{\partial t}|_{t=t_{n}}$ (derivative prediction). From a deep learning perspective, this does not change the model, architecture, or training procedure, aside from the need to compute temporal derivatives from the dataset to use as labels. This can be done numerically, such as with a forward Euler scheme ($\frac{\mathbf{u}(t_{n+1})-\mathbf{u}(t_{n})}{\Delta t}$), central difference scheme ($\frac{\mathbf{u}(t_{n+1})-\mathbf{u}(t_{n-1})}{2\Delta t}$), or higher-order Richardson extrapolations. Additionally, at the endpoints of the dataset a one-sided Richardson extrapolation can be used to obtain accurate estimates of the temporal derivative [46]. By using increasingly higher-order finite-difference schemes, training labels can made increasingly accurate as well. Additionally, the current time is passed into the model through using Fourier features [47] to project the timestep $t_{n}$ to a higher dimension and using Adaptive Layer Normalization. Lastly, depending on the PDE setup, coefficient information can also be provided to the model in a similar manner.

### v2

@@ Abstract
Neural surrogates for partial differential equations (PDEs) have become popular due to their potential to quickly simulate physics. With a few exceptions, neural surrogates generally treat the forward evolution of time-dependent PDEs as a black box by directly predicting the next state. While this is a natural and easy framework for applying neural surrogates, it can be an over-simplified and rigid framework for predicting physics. In this work, we evaluate an alternate framework in which neural solvers predict the temporal derivative and an ODE integrator forwards the solution in time, which has little overhead and is broadly applicable across model architectures and PDEs. We find that by simply changing the training target and introducing numerical integration during inference, neural surrogates can gain accuracy and stability in finely-discretized regimes. Predicting temporal derivatives also allows models to not be constrained to a specific temporal discretization, allowing for flexible time-stepping during inference or training on higher-resolution PDE data. Lastly, we investigate why this framework can be beneficial and in what situations does it work well.
@@ 1 Introduction
Solving equations of this form is of great interest, motivating centuries of research and leading to many modern engineering and scientific advances. Although solutions are analytically intractable, they can be approximated by discretizing the domain, where $t_{n}\in\{t_{0},\ldots,t_{N}\}$, $x_{m}\in\{x_{0},\ldots,x_{M}\}$, and the solution $\mathbf{u}(t_{n},x_{m})$ is discretized in this domain. Within this setup, initial conditions $\mathbf{u}(0,x_{m})$ and boundary conditions $B[\mathbf{u}](t_{n},x_{b}),x_{b}=\{x_{m}:x_{m}\in\partial\Omega\}$ are usually given based on the practical application of the PDE.
@@
Neural Surrogates As a result, deriving and solving an accurate approximation of $\mathbf{u}$ can require significant technical expertise and computational resources. This has motivated a new class of deep learning methods, or neural surrogates, that can learn this update from data, rather than an analytical derivation, and be quickly queried [3, 4]. The fast inference speed of deep learning models makes neural surrogates attractive, however, they lack performance guarantees and the rigorous mathematical theory behind numerical analysis. Regardless, data-driven neural surrogates have shown empirically good performance and the field is rapidly growing to address fundamental challenges in model accuracy, speed, robustness, and practicality.
Many works have proposed neural surrogates based on different deep learning architectures, as well as proposed a diverse set of modifications and architecture changes to tailor networks to physics applications. These include neural surrogates based on neural operators [5, 6], graph neural networks (GNNs) [7, 8], transformers [9, 10, 11], convolutional neural networks (CNNs) [12, 13], or diffusion models [14, 15, 16], each with their different benefits and drawbacks. For example, CNNs are simple to train and empirically show good performance [15], however they can only be applied to regular-grid problems. On the other hand, transformers can be used on arbitrary geometries and perform well; however, they can be costly due to the quadratic complexity of attention, and specialized algorithms have been developed to mitigate this [17, 18].
In addition to advancements in neural network architectures, many works have aimed to improve different aspects of neural surrogates, such as modifying the training strategy or datasets. A few examples of these works include improving the spectral performance of neural surrogates [19], adapting to irregularly discretized data [20, 21], and scaling models to large parameter sizes and across different datasets [22, 23, 24, 25]. The relevant literature is too vast to detail here; however, for a list of representative neural surrogate works and their numerical baselines, we recommend McGreivy and Hakim [26] and for a survey on neural surrogates and operator learning we recommend Kovachki et al. [5]. Within this large and continually expanding body of work, we evaluate an alternate training framework, one that can be applied generally to different architectures and can work together with different training modifications and across datasets.
@@
Contributions Neural surrogates often treat solution updates as a black box and directly predict the next state, which we call state prediction ($\mathbf{u}(t_{n+1})=F_{\theta}(\mathbf{u}(t_{n}))$). State prediction has become the dominant framework for training and querying neural surrogates, as it is the most obvious way to step forward in time, is easy to implement, and allows neural surrogates to take larger steps $\Delta t$. However, we believe that this can be an over-simplification and, inspired by numerical methods, we ask if predicting the temporal derivative or derivative prediction ($\frac{\partial\mathbf{u}}{\partial t}|_{t=t_{n}}=F_{\theta}(\mathbf{u}(t_{n}))$) can improve neural surrogate accuracies. Derivative prediction changes the training objective and inference procedure, and while this framework does not contribute additional training cost, it can introduce additional inference costs. Fortunately, this is the only practical difference between the two frameworks, allowing derivative prediction to be used with any neural surrogate architecture and generally applied to the same range of PDEs, discretizations, and problem setups.
@@
One may ask why it is necessary to investigate alternative neural surrogate training and inference frameworks. We illustrate its differences in Figure 1 and describe its potential benefits here. Our hypothesis is that the main benefit of derivative prediction is that it allows the use of an ODE integrator to evolve the solution based on the predicted temporal derivative, which can be quickly queried at inference time. This effectively decouples the prediction problem, where the model learns the spatial dynamics that create the change in solution values, and the numerical integrator forwards the solution in time based on this learned derivative. In many PDE setups, this can greatly simplify the learning objective; rather than needing to collapse a spatial and temporal update under a single model prediction (state prediction), models can focus on learning just the change in solution values, which can be more important than the actual value of the solutions. Through experimentation, we show that this can result in more accurate model predictions.
@@
Secondly, derivative prediction can be more flexible during inference than state prediction. State prediction fixes the temporal resolution of the solution to the dataset resolution, whereas derivative prediction can generate solutions of arbitrary temporal resolutions by modifying $\Delta t$ in the ODE integrator during inference, which could additionally introduce adaptive step sizing for neural surrogates. One remark is that the temporal resolution can be varied only to a point, since large $\Delta t$ can introduce errors, especially for stiff problems. Another benefit of derivative prediction is that it can take advantage of different integration schemes, such as multi-step, predictor-corrector, or Runge-Kutta methods to improve or adjust solutions during inference. Importantly, under derivative prediction, these modifications during inference do not require re-training the model. We present experimental results demonstrating these additional capabilities.
@@
A subtle but interesting consequence of this is that models trained with derivative prediction can use PDE data more effectively than training in the conventional state prediction framework. A common practice when training neural surrogates is to down-sample PDE data in time, which is usually finely discretized due to numerical solver constraints, and often up to factors of 102-104. This is necessary since predicting the next timestep directly implicitly fixes the temporal resolution of the learned model; training on the full-resolution data would require models to auto-regressively predict a large number of timesteps due to its fine resolution and therefore quickly accumulate error. Unfortunately, downsampling by large factors results in discarding a considerable amount of existing data, but since derivative prediction does not fix the model to a temporal resolution, this can be avoided. In particular, models can be trained on the full-resolution data to obtain highly accurate estimates of the instantaneous derivative, which can be used to take variable steps in time when numerically integrating during inference. We provide results demonstrating this as well.
@@
The benefits of the derivative prediction are not without drawbacks. The main limitation of derivative prediction is that it reintroduces numerical error and discretization constraints back into the neural surrogate framework. Specifically, even with perfect estimates of the instantaneous derivative, numerical integrators can still accumulate error over time with large timesteps $\Delta t$ or stiff PDE dynamics. While this is an issue, we find that in practice, this additional numerical error is very small when compared to model error, especially when using higher-order integrators. Additionally, common timescales that conventional neural surrogates are trained on are generally small enough for ODE integrators to handle, however this is not always the case. This can be seen for steady-state problems where neural surrogates are tasked with predicting a future steady-state solution given solely the initial condition (such as Darcy Flow); derivative prediction will completely fail since the initial derivative has little correlation to the steady-state solution. While we only consider time-dependent PDEs, we seek to understand the limits of neural surrogates when predicting different timescales, both under state prediction and derivative prediction frameworks. In particular, it is known that neural surrogates can take timesteps that are numerically instable (e.g., $CFL>1$), but we seek to understand to what extent this is possible. Lastly, we release the code and datasets for this work here: https://github.com/anthonyzhou-1/temporal_pdes.
@@ 2 Related Work
Neural Surrogate Modifications Initial work on neural surrogates designed various architectures to quickly and accurately approximate PDE solutions, including DeepONet [4], neural operators [3], and physics-informed methods [27]. Although powerful on their own, subsequent works have proposed modifications to these architectures to improve accuracy [28, 29, 18], generalize to different problems [22, 10], or adapt to complex PDE systems [30, 31, 32]. While these modifications are insightful and perform well, they are generally problem- or model-specific, which limits their applicability. To address this, prior work has also considered proposing training modifications or guidelines to improve surrogate accuracies [24]. One such example is the use of data augmentations to improve neural surrogate training and prediction accuracy [33]. In addition, previous work has explored the idea of noising data or unrolled training to improve neural surrogate stability and robustness [20, 22, 34]. Lastly, an interesting set of works seek to train refiner modules alongside neural surrogates to improve solution accuracy [19].
@@
Residual/Derivative Prediction Prior work has also examined learning physics under different training and prediction frameworks. Besides state prediction, the most common training framework is to predict the residual between the input and the target, or residual prediction. Residual prediction can be seen as a special case of derivative prediction, where the residual is a scaled forward Euler approximation of the temporal derivative; however, the temporal resolution is still fixed during inference. Residual prediction is used widely in weather applications [35, 36], and its use in PDE applications can also be found [37]. Pfaff et al. [38] and Sanchez-Gonzalez et al. [39] use residual prediction with GNN-based neural surrogates to approximate a Forward Euler step. Stachenfeld et al. [40] and Wang et al. [41] also use residual prediction with CNN-based neural surrogates, inspired by temporal updates used in numerical methods. Li et al. [32] investigate the effect of residual prediction as temporal resolution decreases, as a part of a larger study in learning chaotic PDE dynamics. Lastly, Sanchez-Gonzalez et al. [42] and Zeng et al. [43] extend these ideas to use an RK2 integrator with a GNN-based architecture, after predicting the derivative using a Hamiltonian framework or with a numerical solver-inspired architecture.
Beyond PDE surrogate modeling, derivative prediction is a common framework the molecular dynamics and machine learned interatomic potential research [44]. Rather than learning the derivative of nodal values, such as velocity, temperature, or pressure, machine learning approaches in the molecular dynamics field seek to predict the energy of a given state. This energy can then be differentiated to obtain the relevant force fields and a loss can be calculated with respect to empirical force fields from ab initio datasets [45, 46, 47].
While the idea of derivative prediction is not novel, we investigate in what situations it is beneficial, why it can help PDE surrogate models, and compare it to state-prediction and other training modifications. In addition, while other works use derivative prediction as a numerical-solver motivated technique to empirically improve surrogate models, we believe that it can be applied more generally to arbitrary models. As such, we investigate derivative prediction as an broadly applicable training framework, rather than a technique applied on an ad-hoc basis.
@@
Hybrid Solvers Another set of more restrictive, but more targeted training frameworks is to use the neural surrogate to approximate quantities within a conventional solver, creating a hybrid solver. For example, when considering the Navier-Stokes equations, hybrid solvers can use a neural surrogate to approximate the convective flux due to its large computational burden and solve the remaining terms or steps numerically [48, 49]. For other equations, spatial derivatives can also be individually approximated using a neural surrogate [50], and more generally, neural surrogates have found use in approximating costly terms within numerical solvers using data [51, 52]. Furthermore, another set of works use a numerical solver to train an inverse neural PDE solver which predicts system inputs or variables from solutions. [53, 54]. While these frameworks benefit from increased performance or capabilities from using a numerical solver, they are also tied to the specific problem to numerical framework (e.g, FDM, FVM, etc.) which can limit their applicability.
@@ 3.1 Training
During training, models parameterized by $F_{\theta}$ are given the current solution $\mathbf{u}(t_{n})$ and timestep $t_{n}$, and are trained to predict either $\mathbf{u}(t_{n+1})$ (state prediction) or $\frac{\partial\mathbf{u}}{\partial t}|_{t=t_{n}}$ (derivative prediction). This results in the following loss function $\mathcal{L}_{\theta}$, for a given neural network parameterized by $F_{\theta}$:
$\mathcal{L}_{\theta}(\mathbf{u}(t_{n}),t_{n},\mathbf{y})=||F_{\theta}(\mathbf{u}(t_{n}),t_{n})-\mathbf{y}||_{2}^{2},\qquad\mathbf{y}=\begin{cases}\mathbf{u}(t_{n+1})&\text{state-prediction}\\ \frac{\partial\mathbf{u}}{\partial t}|_{t=t_{n}}&\text{derivative-prediction}\end{cases}$ (2)
From a deep learning perspective, this does not change the model, architecture, or training procedure, aside from the need to compute temporal derivatives from the dataset to use as labels. This can be done numerically, such as with a forward Euler scheme ($\frac{\mathbf{u}(t_{n+1})-\mathbf{u}(t_{n})}{\Delta t}$), central difference scheme ($\frac{\mathbf{u}(t_{n+1})-\mathbf{u}(t_{n-1})}{2\Delta t}$), or higher-order Richardson extrapolations. Additionally, at the endpoints of the dataset a one-sided Richardson extrapolation can be used to obtain accurate estimates of the temporal derivative [60]. By using increasingly higher-order finite-difference schemes, training labels can made increasingly accurate as well.
In practice, the temporal derivatives for a given dataset can be precomputed and cached for training speed; however, even if derivatives are calculated on the fly during training, we have found the additional overhead to be small. At each training step, a random timestep $t_{n},\>n\in[0,T]$ is uniformly sampled, and the loss $\mathcal{L}_{\theta}(\mathbf{u}(t_{n}),t_{n},\mathbf{y})$ is evaluated. Additional information such as the current timestep or coefficient information is passed into the model using sinusoidal positional embeddings [61, 62] to project the inputs to a higher dimension. Mathematically, given conditioning information $x\in\mathbb{R}$ (this could be $t_{n}$ or coefficients $\nu,\>c,\>\text{etc.}$.), the embedding $\mathbf{e}\in\mathbb{R}^{d}$ is generated by $\gamma:\mathbb{R}\rightarrow\mathbb{R}^{d}$:
$\mathbf{e}=\gamma(x)=[\gamma(x)_{1},\>\gamma(x)_{2},\>\ldots,\>\gamma(x)_{d}]^{T},\quad\text{where}\>\gamma(x)_{i}=\begin{cases}\sin(\frac{x}{10000^{i/d}})&i\text{ is odd}\\ \cos(\frac{x}{10000^{i/d}})&i\text{ is even}\end{cases}$ (3)
After the embeddings are generated, Adaptive Layer Normalization is used to condition model layers on this additional information. Specifically, we consider a neural network defined as a composition of $L$ model layers: $F_{\theta}(x)=\phi_{L}\circ\ldots\circ\phi_{2}\circ\phi_{1}(\mathbf{x})$. Adaptive Layer Normalization uses a separate network at each layer $\psi_{l}(\mathbf{e}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{2d}$ to project embeddings $\mathbf{e}\in\mathbb{R}^{d}$ to a scale and shift parameter $\mathbf{\alpha}\in\mathbb{R}^{d},\>\mathbf{\beta}\in\mathbb{R}^{d}$, where $d$ is the model dimension. At a given layer $l\in[1,2,\ldots,L]$, the hidden activation $\mathbf{h}_{l}\in\mathbb{R}^{d}$ is modulated by:
$\mathbf{h}_{l}^{*}=\phi_{l}(\mathbf{h}_{l-1}),\quad[\mathbf{\alpha},\mathbf{\beta}]^{T}=\psi_{l}(\mathbf{e}),\quad\mathbf{h}_{l}=\mathbf{\alpha}\mathbf{h}_{l}^{*}+\beta$ (4)
A final consideration is that coefficients and timesteps are scaled to be between the range $[0,1]$ to prevent overfitting to their absolute values. Lastly, timestep information is not strictly necessary due to the temporal invariance of dynamical systems, however, including it does not change performance and the results can generalize to architectures that do not provide this information.